OpenStudy (anonymous):

Solve the equation. convert improper fractions to mixed or proper fractions. 2 3/10 + 3 2/10 =

OpenStudy (golikrishna):

apply same here also

ganeshie8 (ganeshie8):

@golikrishna what to apply same here also... you didnt explain him anything in the last question.. you directly giving answers lol.. pls explain the concept also.....

OpenStudy (anonymous):

can SOMEONE please explain how to solve these please ? dangg !

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@SmoothMath : nobod is explaining to me how i solve these kind of problems at alll ! help please ?

OpenStudy (anonymous):

@lgbasallote

OpenStudy (lgbasallote):

\[\huge 2 \frac{3}{10} + 3 \frac{2}{10}\] since it is common denominator there is an easy way to do this \[\huge (2 + 3) \times (\frac{3 + 2}{10})\] remember you can only do this when there'scommon denominators

OpenStudy (anonymous):

i got 2.5 ?

OpenStudy (lgbasallote):

turn that .5 into a fraction

OpenStudy (lgbasallote):

and it'snot 2

OpenStudy (lgbasallote):

what's 2 +3?

OpenStudy (anonymous):

so would it be 5 5/10 ?

OpenStudy (lgbasallote):

yup

OpenStudy (lgbasallote):

you can simplify 5/10 though

OpenStudy (anonymous):

5 1/2 ?

OpenStudy (lgbasallote):

you got it!

OpenStudy (anonymous):

what if they dont have the same denominator for example: 8 1/6 + 2/3 ? would it be 8 3/3 ?

OpenStudy (lgbasallote):

nope...there are 2 ways 1) turn them into improper fractions then take the LCD 2) turn 2/3 into a denominator of 6

OpenStudy (anonymous):

so 2/6 + 49/6 ?

OpenStudy (anonymous):

@lgbasallote ?

OpenStudy (lgbasallote):

what did you do?

OpenStudy (lgbasallote):

oh wait i see

OpenStudy (lgbasallote):

yup that'sright